# INTERVENOR[Uncaptioned image]: Prompt the Coding Ability of Large Language Models with the Interactive Chain of Repairing

### 19 Jan 2024 v3

@@ A.3 Additional Case Studies
Figure 10: Case Studies. We present two cases from CodeError. The chain of repairing clearly identifies the NameError and TypeError in the buggy code and provides the correct modification methods, which are highlighted in red.
@@
Figure 16: Zero-Shot and Few-Shot Prompts for Code Repairing.
Figure 17: Few-Shot CoT Prompts for Code Repairing.
Figure 18: Prompts Used in Self-Refine Model. The instructions for repair plan generation and code repair are shown.
Figure 19: Prompts Used in Error Msgs Model. It conduct code Repairing directly use the code error messages.
A.4 Prompts for Agent Building
@@
We first show the role instructions used by Code Learner and Code Teacher in Figure 13. We utilize specific instructions to ensure that the Code Teacher and Code Learner within INTERVENOR play the correct roles and carry out code inspection and code repair. Then, we provide prompts used in code generation and code repair. In Figures 14 to 15, we show the Zero-Shot, Zero-Shot CoT, Few-Shot, and Few-Shot CoT prompts for code generation tasks. In Figures 16 to 19, we show some code repair prompts, including Zero-Shot, Few-Shot, Few-shot CoT, Self-Refine, and Error Msgs. Self-Refine asks the LLMs to rethink the errors by themselves, while Error Msgs incorporate the code error messages from compilers and directly use them to guide the code repair process.

### 20 Feb 2024 v4

@@ A.4 Additional Case Studies
Figure 10: Case Studies. We present two cases from CodeError. The Chain-of-Repair clearly identifies the NameError and TypeError in the buggy code and provides the correct modification methods, which are highlighted in red.
@@
Figure 16: Zero-Shot and Few-Shot Prompts for Code Repair.
Figure 17: Few-Shot CoT Prompts for Code Repair.
Figure 18: Prompts Used in Self-Refine Model. The instructions for generating code repair plannings and code repair are shown.
Figure 19: Prompts Used in Error Msgs Model. It conduct code repair by directly using the code error messages.
A.5 Prompts for Agent Building
@@
We first show the role-playing instructions used by Code Learner and Code Teacher in Figure 13. We utilize specific instructions to ensure that the Code Teacher and Code Learner within INTERVENOR play the correct roles and carry out code inspection and code repair.
Then, we present the prompts used in code generation and code repair. From Figures 14 to 15, we show the Zero-Shot, Zero-Shot CoT, Few-Shot, and Few-Shot CoT prompts for code generation tasks. From Figures 16 to 19, we show the prompts used in some code repair baseline models, including Zero-Shot, Few-Shot, CoT, Self-Refine, and Error Msgs. Self-Refine asks the LLMs to rethink the errors by themselves, while Error Msgs incorporate the code error messages from compilers and directly use them to guide the code repair process.